• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name slifer

    Log entry time 00:04:33 on June 3,2002

    Entry number 85132

    keyword=swing shift summary

    4:00p.m. waiting for beam
    5:00 Accelerator ready to deliver beam, but we have small target prob(see halog)
    5:20 Beam to hall.
    5:37 kinematic Q1_cdN started. Cross normalization.
    run 1118. 100% DT. kcoda.
    5:42 run 1119 100% DT again. kcoda and rebooted the crates
    run 1120. good run. about 1M events
    6:20 Start kinematic change to Q1_d20
    left arm moved to floor marker 20.50degree, vernier to -50.3==>20.031 degrees
    left arm quadropoles cycled. P0 set to 2.32650 GeV
    6:35 Kinematic change complete. Beam rf cavity problems.
    6:50 Beam to hall. Start 5uA low current run to study tc spectrum. run 1121
    7:30 raised current to 100uA for production running. Beam unstable so we
    requested 70uA instead.
    7:40 run 1122. production @ kinematic q1_d20.
    8:00 MCC informs us of rf cavity problem. Estimates a 15 min downtime.
    8:10 Beam back to hall. Run 1123. Running at 70uA due to rf cavity instability.
    8:27 Beam down. AES will work on injector rf cavity. Estimates 1/2 hour.
    9:20 AES has been called in to work on cavity.

    midnite: Still waiting for beam.